OpenStudy (anonymous):

Find the term c14 in the following sequence: 12, 4, 4/3, 4/9

OpenStudy (anonymous):

here first term and ratio are?

OpenStudy (anonymous):

here first term is 12 common ratio = 1/3 C14 = a*r^(n-1)= 12 * (1/3)^13 = 4/531441